if -2x+7y=-8 and x-6y=10 then what does x-y equal?
Answer by stanbon(75887)

-2x+7y=-8 and x-6y=10 then what does x-y equal?
--------
EQUATIONS:
-2x + 7y = -8
x - 6y = 10
-----------------
x = 6y + 10
Substitute to get:
-2(6y+10) + 7y = -8
-12y - 20 + 7y = -8
-5y = 12
y = -12/5
---------------
Substitute into x = 6y +10 to solve for "x":
x = 6(-12/5) +10
x = (-72/5) + (50/5)
x = -22/5
=================
So x-y = -22/5 - (-12/5) = -10/5 = -2
==================
Cheers,
Stan H.